OpenStudy (anonymous):

please help! two stores had a combined total of 210 teddy bears. after the first store sold 1/2 of it teddy bears and the second store sold 1/3 of its teddy bears, each store had the same number of teddy bears left. how many teddy bears were sold from the two stores combined????

OpenStudy (anonymous):

a+b=210 1/2a=1/3b Simultaneous equations to solve for a or b.

OpenStudy (anonymous):

b+2/3b=210

OpenStudy (anonymous):

so 5/3b=210, what is b

OpenStudy (anonymous):

and then 210-b= a

OpenStudy (anonymous):

Then you can find half of one and a third of the other

OpenStudy (anonymous):

Close the question if you are done

OpenStudy (anonymous):

After the second store sold 1/3 of its stock, it had 2/3 left. It is at this point that their stock levels were equal. The second equation in the set up should be (1/2)a=(2/3)b.